Two telephone poles, 30 ft and 39 ft high, respectively, are 60 ft apart. How long is the wire from the top of one pole to the top of the other pole?
Answer by checkley77(12844)

Here you have a right triangl with one side=60, the other side is 39-30=9.
60^2+9^2=h^2
3600+81=h^2
3681=h^2
h=sqrt3681
h=60.67 feet for the length of the wire.